Use the data in the basket of goods table to answer the five questions below. Please show your work.

Basket of Goods

Good

Quantity

Price in base year

Price in current year

Pepperoni pizza

15

$15.00

$17.50

Bottles of soda

30

$1.25

$1.50

Movie Tickets

12

$8.00

$10.00

4. The CPI for the current year is

 

 

5. The percentage change in the CPI from the base year to the current year was

 

 

 

 



1
Expert's answer
2021-02-24T15:14:08-0500

"CPI for the basket=\frac{Cost of market basket t}{Cost of market basket tb}\times100"

"CPI for the basket=\frac{Cost of market basket t}{Cost of market basket tb}\times100=\frac{15\times17.5+30\times1.5+12\times10}{15\times15+30\times1.25+12\times8}\times100=\frac{262.5+45+120}{225+37.5+96}=\frac{427.5}{358.5}\times100=119.25"



"ThepercentagechangeintheCPIfromthebaseyeartothecurrentyear=\\frac{(Costcurrentyear\u2212Costbaseyear)}{Costbaseyear}\u200b\u00d7100=\\frac{427.5-358.5}{358.5}\\times100=19.25"